Grade 12th pass9 grade science

When brakes are applied the velocity of a car changes from 30 ms−1to 10 ms−1 in 5 s. The acceleration produced in it is ____ ms−2.

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
dear student
your answer is
RETARDATION =30-0/5=6ms-2.  acceleration =v-u/t     acceleration =v-u/t
thanks